# Numerical investigation of buoyancy-aided mixed convective flow past a square cylinder inclined at 45 degrees

### v3

@@ III.1 Instantaneous Flow
The forced convective component of the flow causes shear layer separation at the transverse corners of the cylinder, causing a deficit of momentum in the wake. On the other hand, the natural convective component causes jet-like behavior downstream, competing with the momentum deficit caused by the wake, with the respective vorticities signed oppositely to each other. The opposing effects of these two flow components lead to a vorticity inversion. This inversion does not occur for Ri $=0.0$ due to the absence of buoyancy at this Ri. For positive Ri, vorticity inversion occurs downstream at an inception point, which comes further upstream as Ri is increased (see Figure 15) due to the increased buoyancy (see Figures 3 and 9).
Isotherms also reveal the presence of vortex shedding in the near-field for Ri $=0.0$ and the steady far-field, as confirmed by examining temperature signals. At Ri $=0.2$, the undulation of shear layers extends to Y $=65$, and the complete flow field is unsteady up to Ri $=0.6$. The shape of the isotherm structures continues to change when Ri is increased due to the increased strength of the buoyancy force and the constant strength of the inertial force. With a further increase in Ri, the temperature in the near-field becomes steady, while the far-field reveals the presence of structures similar to those seen in plumes, as shown by Kimura and Bejan [31]. The near-field plots of temperature show similar behavior as the vorticity contours (see Figure 7). The undulation of the isotherms becomes significantly lower with increasing Ri, and for Ri $\geq$ 0.7, the structures reveal a steady flow and thermal field in the near-field.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
Figure 6: Instantaneous temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
Figure 7: Instantaneous temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Y $=11.5$
(b) Y $=73.5$
(c) Y $=11.5$
(d) Y $=73.5$
@@
Figure 8: Streamwise (a, b) and transverse (c, d) velocity signals in the near- and far-field for Ri $=0.6$ and Ri $=0.7$, along with (e) drag and lift coefficient signals. The near-field signals reveal the suppression of vortex shedding, and the far-field signals show the far-field unsteadiness. In contrast, the force coefficient signals characterize the near-field flow excellently.
To confirm the behavior shown by the flow field in the contours, we have plotted time signals of the streamwise and transverse velocity in the near-field (Y $=11.5$) and far-field (Y $=73.5$), which are shown in Figure 8 for Ri $=0.6$ and Ri $=0.7$. For Ri $=0.6$, streamwise and transverse velocity plots in the near-field reveal unsteady, periodic behavior, and for Ri $=0.7$, a steady behavior without any periodicity is observed (see Figures 8(a) and 8(c)). Far-field streamwise and transverse velocity plots show the opposite behavior, with negligible unsteadiness in the case of Ri $=0.6$ but a large-scale unsteadiness for Ri $=0.7$ associated with the onset of the far-field plume-like unsteadiness (see Figures 8(b) and 8(d)).
The plots of lift and drag coefficient (see Figure 8(e)) show similarities with the near-field velocity signals, suggesting that the near-field flow is unsteady for Ri $=0.6$ and steady for Ri $=0.7$, supporting the suppression of vortex shedding in the near-field. This makes the force coefficients excellent predictors of the nature of the near-field flow.
Table 6: Variation of Strouhal number (St) with Richardson number (Ri) in the near-field (Y $=11.5$) and far-field (Y $=73.5$). The transverse velocity signal along the centerline (X $=0$) has been used to calculate the Strouhal number.
Richardson number (Ri) Strouhal number (St)
Near-field (Y $=11.5$, X $=0$) Far-field (Y $=73.5$, X $=0$)
@@
Table 6 shows the variation of Strouhal number (St) with Ri. The transverse velocity signal along the centerline (X $=0$) is used to compute the Strouhal number at two different streamwise locations, Y $=11.5$ and Y $=73.5$, which fall in the near-field and far-field, respectively. St shows negligible variation between the near-field and far-field for Ri $<$ 0.7. Above the critical Ri, the near-field flow has no periodicity, and the corresponding St is zero. The flow in the wake is accelerated due to the buoyancy force, which is more significant at elevated Richardson numbers, resulting in increased Strouhal number with increasing Ri.
@@ III.2 Time-Averaged Flow
Time-averaged vorticity and temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$ are presented in Figures 9 and 11 in the full domain and Figures 10 and 12 in the vicinity of the cylinder, respectively. The time-averaged vorticity contours in the complete domain (see Figure 9) reveal that at Ri $=0.0$, the strength of the vorticity is high in the near-field but diminishes in the far-field. In contrast, for Ri $>$ 0.0, the strength of the vorticity is weak in the near-field but grows on moving downstream, showing the opposite nature of the forced convective vorticity and the natural convective vorticity. The forced convective vorticity is introduced due to the shear layer separation at the transverse/lateral corners of the cylinder, while the natural convective vorticity is introduced due to buoyancy, which causes a jet-like effect in the wake. The interplay of these opposing vorticities is the cause of many physical phenomena that we have observed in this study and will be discussed in detail later. Beyond Ri $=0.7$, the strength of the vorticity is significantly higher than that of the cases below Ri $=0.7$, and the oppositely-signed vortices remain attached up to the far-field, which was not the case at lower Ri. Further, the two vortices are found to separate suddenly in the far-field at a point (around Y $=65$) which also coincides with the point of onset of the far-field unsteadiness. All positive Ri cases display vorticity inversion, as seen in the instantaneous contours earlier. In the near-field plot of vorticity, the width between the shear layers is seen to reduce with increasing Ri, similar to the instantaneous case. Beyond Ri $=0.7$, the inversion can be seen in the near-field, and the vortices form a steady recirculation bubble, whose length decreases with increasing Ri. These cases resemble the instantaneous plots as the flow is steady in the near-field. Time-averaged temperature contours (see Figure 11) show a widening of the plume in the near and intermediate fields for Ri $<$ 0.7. Beyond Ri=0.7, the plume is narrow in the near-field, and there is a sharp increase in the plume width in the far-field, associated with the onset of the far-field unsteadiness. Such behavior is also observed in the near-field temperature contours (see Figure 12), where the plume is wide in the near-field for low Ri, and beyond Ri $=0.7$, the plume is narrow.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 9: Time-averaged vorticity contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 10: Time-averaged vorticity contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 11: Time-averaged temperature contours for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
(a) Ri $=0.0$
(b) Ri $=0.2$
(c) Ri $=0.4$
(d) Ri $=0.6$
(e) Ri $=0.7$
(f) Ri $=1.0$
@@
Figure 12: Time-averaged temperature contours (in the cylinder vicinity) for Re $=100$ and $0.0\leq$ Ri $\leq 1.0$: (a) Ri $=0.0$, (b) Ri $=0.2$, (c) Ri $=0.4$, (d) Ri $=0.6$, (e) Ri $=0.7$, (f) Ri $=1.0$.
@@
Figure 13 shows the centerline streamwise velocity and centerline temperature along the streamwise length of the domain for different Ri. Figure 13(a) reveals the buoyancy-driven jetting effect, which becomes more pronounced as one moves further downstream. The average streamwise velocity ($V_{avg}$) at any particular downstream location Y is equal to the free-stream velocity ($V_{\infty}$ = 1.0, in the present case). It is to be noted that the wake has a momentum deficit when the centerline streamwise velocity is lower than $V_{avg}$ and becomes a wake with an excess momentum if the centerline velocity exceeds $V_{avg}$ at that location. There is a deficit of momentum in the immediate wake of the cylinder due to shear layer separation at the front lateral corners. However, the presence of the natural convective flow counteracts this deficit and accelerates the flow behind the cylinder, leading to a faster recovery and, eventually, an excess of momentum beyond the cylinder exhibiting jet-like characteristics far downstream, while the flow immediately downstream of the cylinder shows a wake-like behavior in the near-field. For the purpose of visualizing the location of momentum recovery and subsequent excess, $\overline{V}$ = 1.0 has been plotted in Figure 13(a). The centerline velocity variation reveals that the momentum deficit persists in the forced convection case throughout the domain despite the recovery of some momentum later downstream. For positive Ri, the momentum recovery is aided by the buoyancy-driven jet, and the recovery distance decreases with increasing Ri, occurring at Y $=20$ for Ri $=0.2$ and around Y $=5$ for Ri $\geq$ 0.6. Further downstream, for Ri $\geq$ 0.7, there is a decrease in velocity in the far-field, which occurs around the same distance as the onset of the far-field unsteadiness. This phenomenon is attributed to the widening of the wake caused by the unsteadiness of the flow beyond this location. The inset in the same graph shows the centerline velocity in the immediate wake of the cylinder (0.5 $\leq$ Y $\leq$ 1.75), providing an insight into the recirculation length. Starting from the base of the cylinder (Y $=0.5$), there is a negative streamwise velocity due to the recirculation of flow as it turns towards the cylinder (in the time-averaged sense). However, the velocity quickly becomes positive after the recirculation region. The recirculation length shows a nonlinear variation, with Ri $=1$ and Ri $=0.2$ having the shortest recirculation length, followed by Ri $=0.0$, 0.6 and 0.7, respectively. Figure 13(b) shows the centreline variation of temperature with streamwise distance, revealing a reducing trend of temperature with downstream distance. The centerline temperature is highest for higher Ri close to the cylinder; however, in the far-field, the behavior is the opposite, and higher Ri have lower temperatures, albeit the difference is marginal in the far-field.
@@
Figure 13: Variation of (a) centerline streamwise velocity and (b) centerline temperature.
@@
Nusselt number is a function of the spatial coordinates as well as time. Therefore, one can average the Nusselt number in space, time, or both. $\overline{Nu}_{local}$ refers to the time-averaged local Nusselt number at a particular location on the cylinder, while $\overline{Nu}_{surface}$ refers to the temporally and spatially averaged Nusselt number for a particular surface on the cylinder, and $\overline{Nu}$ refers to the temporally and spatially averaged Nusselt number for the entire cylinder. Figure 14 shows the variation of time-averaged Nusselt number for different cylinder surfaces and local Nusselt number along the cylinder surface for various Ri, along with the nomenclature of the cylinder surfaces (AB, BC, CD, and DA). For flow-facing surfaces (AB and DA), the mean Nusselt number increases with increasing Ri, whereas for wake-facing surfaces (BC and CD), the mean Nusselt number decreases with increasing Ri. There is a sharp decrease in the Nusselt number at Ri $=0.7$ for the wake-facing surfaces and, subsequently, in the Nusselt number for the full cylinder. The shedding phenomenon helps in higher heat transfer due to increasing momentum exchange between the thermal boundary layers at various walls and the separating shear layers. When vortex shedding is suppressed, the momentum transport is inhibited as the mutual interaction is no longer active. Therefore, the Nusselt number sharply decreases for the wake-facing surfaces at Ri $=0.7$. The Nusselt number for the surfaces AB and DA, as well as that for surfaces BC and CD, are equal due to the symmetric nature of the time-averaged flow about the centerline. The local Nusselt number is the highest for the vertex A, which directly faces the forced flow, and it is considerably lower for the wake-facing surfaces compared to the flow-facing surfaces due to the recirculation region having very weak streamwise velocities, causing decreased heat transfer from the wake-facing surfaces. The Nusselt number reaches a local maxima at the lateral corners of the cylinder (B and D), due to the acceleration of flow at these corners.
@@
Richardson number (Ri) Drag coefficient ($C_{D}$) RMS lift coefficient ($C_{L,\text{rms}}$) Mean Nusselt number ($\overline{Nu}$)
@@
The variation of the integral parameters with Ri is listed in Table 7. $\overline{C}_{D}$ and $\overline{Nu}$ show an increase with increasing Ri. The buoyancy in the vicinity of the cylinder increases both pressure drag, and skin friction drag on the cylinder. On the other hand, as Ri increases, RMS quantities decrease as expected. All quantities presented in Table 7 are near-field quantities, and hence, when the near-field is stabilized, these values become constant with no periodicity. Hence, RMS quantities fall to zero beyond the critical Ri. The RMS Nusselt number has not been reported as it is extremely small for all Ri. The mean Nusselt number presents an interesting variation. Below Ri $=0.7$, the Nusselt number increases monotonically, decreasing to a local minimum value at Ri $=0.7$ and subsequently increasing for Ri $>$ 0.7. Below 0.7, flow acceleration due to buoyancy forces causes thinning of the thermal boundary layer, leading to higher Nusselt numbers for the entire cylinder. There is a greater increase in the Nusselt number of the flow-facing surfaces compared to the decrease in the Nusselt number of the wake-facing surfaces, leading to an overall increase in the Nusselt number of the cylinder. However, at Ri $=0.7$, as discussed earlier, the Nu of the wake-facing surfaces sharply decreases due to the inhibition of vortex shedding, and this decrease is greater than the increase in the Nusselt number of flow-facing surfaces, leading to an overall decrease in the Nusselt number of the cylinder. Beyond Ri $=0.7$, an increase in cylinder Nu is seen, similar to the trend below Ri $=0.7$.

### v4

@@ III.1 Instantaneous Flow
The forced convective component of the flow causes shear layer separation at the transverse corners of the cylinder, causing a deficit of momentum in the wake. On the other hand, the natural convective component causes jet-like behavior downstream, competing with the momentum deficit caused by the wake, with the respective vorticities signed oppositely to each other. The opposing effects of these two flow components lead to a vorticity inversion. This inversion does not occur for Ri =0.0=0.0 due to the absence of buoyancy at this Ri. For positive Ri, vorticity inversion occurs downstream at an inception point, which comes further upstream as Ri is increased (see Figure 15) due to the increased buoyancy (see Figures 3 and 9).
Isotherms also reveal the presence of vortex shedding in the near-field for Ri =0.0=0.0 and the steady far-field, as confirmed by examining temperature signals. At Ri =0.2=0.2, the undulation of shear layers extends to Y =65=65, and the complete flow field is unsteady up to Ri =0.6=0.6. The shape of the isotherm structures continues to change when Ri is increased due to the increased strength of the buoyancy force and the constant strength of the inertial force. With a further increase in Ri, the temperature in the near-field becomes steady, while the far-field reveals the presence of structures similar to those seen in plumes, as shown by Kimura and Bejan [42]. The near-field plots of temperature show similar behavior as the vorticity contours (see Figure 7). The undulation of the isotherms becomes significantly lower with increasing Ri, and for Ri $\geq$ 0.7, the structures reveal a steady flow and thermal field in the near-field.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
Figure 6: Instantaneous temperature contours for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\theta_{min},\theta_{max},\Delta\theta)\equiv(0,0.1,0.001)$.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
Figure 7: Instantaneous temperature contours (in the cylinder vicinity) for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\theta_{min},\theta_{max},\Delta\theta)\equiv(0,0.1,0.001)$.
(a) Y =11.5=11.5
(b) Y =73.5=73.5
(c) Y =11.5=11.5
(d) Y =73.5=73.5
@@
Figure 8: Streamwise (a, b) and transverse (c, d) velocity signals in the near- and far-field for Ri =0.6=0.6 and Ri =0.7=0.7, along with (e) drag and lift coefficient signals. The near-field signals reveal the suppression of vortex shedding, and the far-field signals show the far-field unsteadiness. In contrast, the force coefficient signals characterize the near-field flow excellently.
To confirm the behavior shown by the flow field in the contours, we have plotted time signals of the streamwise and transverse velocity in the near-field (Y =11.5=11.5) and far-field (Y =73.5=73.5), which are shown in Figure 8 for Ri =0.6=0.6 and Ri =0.7=0.7. For Ri =0.6=0.6, streamwise and transverse velocity plots in the near-field reveal unsteady, periodic behavior, and for Ri =0.7=0.7, a steady behavior without any periodicity is observed (see Figures 8(a) and 8(c)). Far-field streamwise and transverse velocity plots show the opposite behavior, with negligible unsteadiness in the case of Ri =0.6=0.6 but a large-scale unsteadiness for Ri =0.7=0.7 associated with the onset of the far-field plume-like unsteadiness (see Figures 8(b) and 8(d)).
The plots of lift and drag coefficient (see Figure 8(e)) show similarities with the near-field velocity signals, suggesting that the near-field flow is unsteady for Ri =0.6=0.6 and steady for Ri =0.7=0.7, supporting the suppression of vortex shedding in the near-field. This makes the force coefficients excellent predictors of the nature of the near-field flow.
Table 6: Variation of Strouhal number (St) with Richardson number (Ri) in the near-field (Y =11.5=11.5) and far-field (Y =73.5=73.5). The transverse velocity signal along the centerline (X =0=0) has been used to calculate the Strouhal number.
Richardson
number (Ri)
Strouhal number (St)
Near-field (Y =11.5=11.5, X =0=0) Far-field (Y =73.5=73.5, X =0=0)
@@
Table 6 shows the variation of Strouhal number (St) with Ri. The transverse velocity signal along the centerline (X =0=0) is used to compute the Strouhal number at two different streamwise locations, Y =11.5=11.5 and Y =73.5=73.5, which fall in the near-field and far-field, respectively. St shows negligible variation between the near-field and far-field for Ri << 0.7. Above the critical Ri, the near-field flow has no periodicity, and the corresponding St is zero. The flow in the wake is accelerated due to the buoyancy force, which is more significant at elevated Richardson numbers, resulting in increased Strouhal number with increasing Ri.
@@ III.2 Time-Averaged Flow
Time-averaged vorticity and temperature contours for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$ are presented in Figures 9 and 11 in the full domain and Figures 10 and 12 in the vicinity of the cylinder, respectively. The time-averaged vorticity contours in the complete domain (see Figure 9) reveal that at Ri =0.0=0.0, the strength of the vorticity is high in the near-field but diminishes in the far-field. In contrast, for Ri >> 0.0, the strength of the vorticity is weak in the near-field but grows on moving downstream, showing the opposite nature of the forced convective vorticity and the natural convective vorticity. The forced convective vorticity is introduced due to the shear layer separation at the transverse/lateral corners of the cylinder, while the natural convective vorticity is introduced due to buoyancy, which causes a jet-like effect in the wake. The interplay of these opposing vorticities is the cause of many physical phenomena that we have observed in this study and will be discussed in detail later. Beyond Ri =0.7=0.7, the strength of the vorticity is significantly higher than that of the cases below Ri =0.7=0.7, and the oppositely-signed vortices remain attached up to the far-field, which was not the case at lower Ri. Further, the two vortices are found to separate suddenly in the far-field at a point (around Y =65=65) which also coincides with the point of onset of the far-field unsteadiness. All positive Ri cases display vorticity inversion, as seen in the instantaneous contours earlier. In the near-field plot of vorticity, the width between the shear layers is seen to reduce with increasing Ri, similar to the instantaneous case. Beyond Ri =0.7=0.7, the inversion can be seen in the near-field, and the vortices form a steady recirculation bubble, whose length decreases with increasing Ri. These cases resemble the instantaneous plots as the flow is steady in the near-field. Time-averaged temperature contours (see Figure 11) show a widening of the plume in the near and intermediate fields for Ri << 0.7. Beyond Ri=0.7, the plume is narrow in the near-field, and there is a sharp increase in the plume width in the far-field, associated with the onset of the far-field unsteadiness. Such behavior is also observed in the near-field temperature contours (see Figure 12), where the plume is wide in the near-field for low Ri, and beyond Ri =0.7=0.7, the plume is narrow.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 9: Time-averaged vorticity contours for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\overline{\omega}_{min},\overline{\omega}_{max},\Delta\overline{\omega})\equiv(-0.5,0.5,0.01)$.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 10: Time-averaged vorticity contours (in the cylinder vicinity) for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\overline{\omega}_{min},\overline{\omega}_{max},\Delta\overline{\omega})\equiv(-4,4,0.12)$.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 11: Time-averaged temperature contours for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\overline{\theta}_{min},\overline{\theta}_{max},\Delta\overline{\theta})\equiv(0,0.1,0.001)$.
(a) Ri =0.0=0.0
(b) Ri =0.2=0.2
(c) Ri =0.4=0.4
(d) Ri =0.6=0.6
(e) Ri =0.7=0.7
(f) Ri =1.0=1.0
@@
Figure 12: Time-averaged temperature contours (in the cylinder vicinity) for Re =100=100 and $0.0\leq$ Ri $\leq 1.0$: (a) Ri =0.0=0.0, (b) Ri =0.2=0.2, (c) Ri =0.4=0.4, (d) Ri =0.6=0.6, (e) Ri =0.7=0.7, (f) Ri =1.0=1.0; ($\overline{\theta}_{min},\overline{\theta}_{max},\Delta\overline{\theta})\equiv(0.05,1,0.04)$.
@@
Figure 13 shows the centerline streamwise velocity and centerline temperature along the streamwise length of the domain for different Ri. Figure 13(a) reveals the buoyancy-driven jetting effect, which becomes more pronounced as one moves further downstream. The average streamwise velocity ($V_{avg}$) at any particular downstream location Y is equal to the free-stream velocity ($V_{\infty}$ = 1.0, in the present case). It is to be noted that the wake has a momentum deficit when the centerline streamwise velocity is lower than $V_{avg}$ and becomes a wake with an excess momentum if the centerline velocity exceeds $V_{avg}$ at that location. There is a deficit of momentum in the immediate wake of the cylinder due to shear layer separation at the front lateral corners. However, the presence of the natural convective flow counteracts this deficit and accelerates the flow behind the cylinder, leading to a faster recovery and, eventually, an excess of momentum beyond the cylinder exhibiting jet-like characteristics far downstream, while the flow immediately downstream of the cylinder shows a wake-like behavior in the near-field. For the purpose of visualizing the location of momentum recovery and subsequent excess, $\overline{V}$ = 1.0 has been plotted in Figure 13(a). The centerline velocity variation reveals that the momentum deficit persists in the forced convection case throughout the domain despite the recovery of some momentum later downstream. For positive Ri, the momentum recovery is aided by the buoyancy-driven jet, and the recovery distance decreases with increasing Ri, occurring at Y =20=20 for Ri =0.2=0.2 and around Y =5=5 for Ri $\geq$ 0.6. Further downstream, for Ri $\geq$ 0.7, there is a decrease in velocity in the far-field, which occurs around the same distance as the onset of the far-field unsteadiness. This phenomenon is attributed to the widening of the wake caused by the unsteadiness of the flow beyond this location. The inset in the same graph shows the centerline velocity in the immediate wake of the cylinder (0.5 $\leq$ Y $\leq$ 1.75), providing an insight into the recirculation length. Starting from the base of the cylinder (Y =0.5=0.5), there is a negative streamwise velocity due to the recirculation of flow as it turns towards the cylinder (in the time-averaged sense). However, the velocity quickly becomes positive after the recirculation region. The recirculation length shows a nonlinear variation, with Ri =1=1 and Ri =0.2=0.2 having the shortest recirculation length, followed by Ri =0.0=0.0, 0.6 and 0.7, respectively. Figure 13(b) shows the centreline variation of temperature with streamwise distance, revealing a reducing trend of temperature with downstream distance. The centerline temperature is highest for higher Ri close to the cylinder; however, in the far-field, the behavior is the opposite, and higher Ri have lower temperatures, albeit the difference is marginal in the far-field.
@@
Figure 13: Variation of (a) centerline streamwise velocity and (b) centerline temperature with streamwise distance.
@@
Nusselt number is a function of the spatial coordinates as well as time. Therefore, one can average the Nusselt number in space, time, or both. $\overline{Nu}_{local}$ refers to the time-averaged local Nusselt number at a particular location on the cylinder, while $\overline{Nu}_{surface}$ refers to the temporally and spatially averaged Nusselt number for a particular surface on the cylinder, and $\overline{Nu}$ refers to the temporally and spatially averaged Nusselt number for the entire cylinder. Figure 14 shows the variation of time-averaged Nusselt number for different cylinder surfaces and local Nusselt number along the cylinder surface for various Ri, along with the nomenclature of the cylinder surfaces (AB, BC, CD, and DA). For flow-facing surfaces (AB and DA), the mean Nusselt number increases with increasing Ri, whereas for wake-facing surfaces (BC and CD), the mean Nusselt number decreases with increasing Ri. There is a sharp decrease in the Nusselt number at Ri =0.7=0.7 for the wake-facing surfaces and, subsequently, in the Nusselt number for the full cylinder. The shedding phenomenon helps in higher heat transfer due to increasing momentum exchange between the thermal boundary layers at various walls and the separating shear layers. When vortex shedding is suppressed, the momentum transport is inhibited as the mutual interaction is no longer active. Therefore, the Nusselt number sharply decreases for the wake-facing surfaces at Ri =0.7=0.7. The Nusselt number for the surfaces AB and DA, as well as that for surfaces BC and CD, are equal due to the symmetric nature of the time-averaged flow about the centerline. The local Nusselt number is the highest for the vertex A, which directly faces the forced flow, and it is considerably lower for the wake-facing surfaces compared to the flow-facing surfaces due to the recirculation region having very weak streamwise velocities, causing decreased heat transfer from the wake-facing surfaces. The Nusselt number reaches a local maxima at the lateral corners of the cylinder (B and D), due to the acceleration of flow at these corners.
@@
Richardson
number (Ri)
Drag coefficient ($C_{D}$)
RMS lift coefficient
($C_{L,\text{rms}}$)
Mean Nusselt number
($\overline{Nu}$)
@@
The variation of the integral parameters with Ri is listed in Table 7. $\overline{C}_{D}$ and $\overline{Nu}$ show an increase with increasing Ri. The buoyancy in the vicinity of the cylinder increases both pressure drag, and skin friction drag on the cylinder. On the other hand, as Ri increases, RMS quantities decrease as expected. All quantities presented in Table 7 are near-field quantities, and hence, when the near-field is stabilized, these values become constant with no periodicity. Hence, RMS quantities fall to zero beyond the critical Ri. The RMS Nusselt number has not been reported as it is extremely small for all Ri. The mean Nusselt number presents an interesting variation. Below Ri =0.7=0.7, the Nusselt number increases monotonically, decreasing to a local minimum value at Ri =0.7=0.7 and subsequently increasing for Ri >> 0.7. Below 0.7, flow acceleration due to buoyancy forces causes thinning of the thermal boundary layer, leading to higher Nusselt numbers for the entire cylinder. There is a greater increase in the Nusselt number of the flow-facing surfaces compared to the decrease in the Nusselt number of the wake-facing surfaces, leading to an overall increase in the Nusselt number of the cylinder. However, at Ri =0.7=0.7, as discussed earlier, the Nu of the wake-facing surfaces sharply decreases due to the inhibition of vortex shedding, and this decrease is greater than the increase in the Nusselt number of flow-facing surfaces, leading to an overall decrease in the Nusselt number of the cylinder. Beyond Ri =0.7=0.7, an increase in cylinder Nu is seen, similar to the trend below Ri =0.7=0.7.
The critical Ri for the transition to steady flow for both square and circular cylinders is 0.15, as reported by Sharma and Eswaran [23], and Chang and Sa [24]. As reported by [36], the corresponding value for an elliptical cylinder with an aspect ratio (AR) of 2 is between 0.75 and 1.0, while for an AR = 4, no vortex shedding suppression is observed for the range of Ri (0.0-1.25) considered in their study. However, the critical Ri for our inclined square cylinder in the present study is between 0.6 and 0.7, which is significantly greater than that reported for square and circular cylinder cases. This large variation may be attributed to the (i) thickness and vorticity content of the separating shear layer and/or (ii) afterbody effect, that is, the part of the obstacle beyond the separation of the flow from it. The shape of the obstacle and the corresponding afterbody are expected to affect the transition sequence in mixed convective flows past obstacles. The exact details of the effect require further investigation.
A comparison of integral parameters among obstacles of different shapes (square [23], circular [24], and elliptical cylinders [36]) against the diamond-shaped cylinder of the present study is given in Table 8. It is seen that the drag coefficient at Ri =0.0=0.0 is highest in the case of an elliptical cylinder with AR = 4, followed by a diamond cylinder, an elliptical cylinder with AR = 2, a square cylinder, and finally, a circular cylinder. A similar trend is observed at higher Ri with minor changes. Similarly, the elliptical cylinders reveal the highest Nusselt number, followed by diamond and square cylinders at all Ri. The Strouhal number is approximately the same for elliptical and diamond-shaped cylinders, while it is considerably lower for square and circular cylinders.
Table 8: Variation of drag coefficient, Nusselt number, and Strouhal number with Ri for different cylinder shapes.
Cylinder
shape
Richardson
number (Ri)
Mean drag
coefficient ($\overline{C}_{D}$)
Mean Nusselt
number ($\overline{Nu}$)
Strouhal
number (St)
Square ([23]) 0.0 1.57 4.09 0.147
Circular ([24]) 0.0 1.22 5.22 0.154
Elliptical (AR = 2) ([36]) 0.0 1.77 5.93 0.183
Elliptical (AR = 4) ([36]) 0.0 2.10 6.26 0.176
Diamond (present) 0.0 1.81 5.25 0.180
Square ([23]) 0.5 2.07 4.55 Steady
Elliptical (AR = 2) ([36]) 0.5 2.14 6.17 0.218
Elliptical (AR = 4) ([36]) 0.5 2.82 6.65 0.224
Diamond (present) 0.5 2.11 5.40 0.212
Square ([23]) 1.0 2.63 4.87 Steady
Elliptical (AR = 2) ([36]) 1.0 2.43 6.15 Steady
Elliptical (AR = 4) ([36]) 1.0 3.25 6.84 Steady
Diamond (present) 1.0 2.44 5.47 Steady